OpenStudy (anonymous):

If 17% of the listeners of a radio station listen to it while they are at work, what is the approximate standard deviation of the sampling distribution of the proportion for samples of size 91? A. 7.8% B. 3.9% C. 1.8% D. 0.4%

jimthompson5910 (jim_thompson5910):

Use the formula \[\Large \sqrt{\frac{p(1-p)}{n}}\]

jimthompson5910 (jim_thompson5910):

In this case, p = 0.17 n = 91

OpenStudy (anonymous):

ok i did is D?

jimthompson5910 (jim_thompson5910):

what result did you get?

jimthompson5910 (jim_thompson5910):

numeric result

OpenStudy (anonymous):

it was 0.039and bunch of other numbers

jimthompson5910 (jim_thompson5910):

convert that to a percentage (multiply the result by 100)

jimthompson5910 (jim_thompson5910):

ex: 0.17 = 0.17*(100%) = 17%

OpenStudy (anonymous):

ok so it would be 3.9377% and the closest answers is 4% so D? :)

jimthompson5910 (jim_thompson5910):

read D again carefully

OpenStudy (anonymous):

Oh wow im an idiot sorry its B right?

jimthompson5910 (jim_thompson5910):

yes B

OpenStudy (anonymous):

thank you so much :D

jimthompson5910 (jim_thompson5910):

np